# Classical Apéry-Like Series and Their Cyclotomic Parametric Analogues via Contour Integration

### v1

@@ 1 Introduction
The primary objective of this paper is to investigate the following cyclotomic parametric Apéry-like series via the method of contour integration:
@@
where $a\in\mathbb{C}\setminus\mathbb{Z},\ q\in\mathbb{N}$ and $z$ is an arbitrary root of unity with $(q,z)\neq(1,1)$. In particular, when $a$ is a half-integer in (1.2) and $2n+2a$ is a non-positive integer, the reciprocal of the parametric central binomial coefficient, $\binom{2n+2a}{n+a}^{-1}$, is interpreted as zero by convention. Here the generalized binomial coefficient $\binom{a}{b}$ is denoted by
@@ 2 Main Theorems and Corollaries
The second cyclotomic Apéry-like series in equation (2.1) above can be expressed in terms of multiple polylogarithms and $\log(2)$, as derived from the following proposition (see [34, Theorem 3.8]).
@@
In [34, Theorem 3.8], the formula omits two absolute value symbol; specifically $\log^{p-1-k}(x)$ should be written as $\log^{p-1-k}(|x|)$, and $\log^{k-j-l}(1-G_{m}(x)^{-1})$ should be written as $\log^{k-j-l}(|1-G_{m}^{-1}(x)|)$. This adjustment is necessary because it can be proven that for $x\in[-(m-1)^{m-1}/m^{m},(m-1)^{m-1}/m^{m}]$, the value of $G_{m}(x)\ (m\geq 2)$ lies within the interval $(1/2,2]$. In Section 5, we will employ a method similar to that used in the proof of [34, Theorem 3.8] to derive an alternative expression for equation (2.2). This approach will allow us to establish several identities involving multiple polylogarithm functions. Certainly, a direct integration approach should yield an alternative expression for equation (2.2). As can be readily observed from (2.9), the series on the left-hand side of (2.2) can be transformed into the following integral:
@@
Based on this integral, following a process similar to the proof of [32, Theorem 2.2], the desired result should be obtainable.
@@
which converges if $|x_{1}\cdots x_{j}|<1$ for all $j=1,\dotsc,r$. In particular, if $x_{1}=x,x_{2}=\cdots=x_{r}=1$, then $\operatorname{Li}_{k_{1},\ldots,k_{r}}(x,\{1\}_{r-1})$ is the single-variable multiple polylogarithm function $\operatorname{Li}_{k_{1},\ldots,k_{r}}(x)$. Clearly, $\operatorname{Li}_{1}(x)=-\log(1-x)\ (x\in[-1,1))$. The (2.11) can be analytically continued to a multi-valued meromorphic function on $\mathbb{C}^{r}$ (see [24, 38, 39]). In [23, Lemma 2.2], Kaneko and Tsumura provided a detailed explanation of how to analytically continue the single-variable multiple polylogarithm to the domain $\mathbb{C}\setminus[1,+\infty)$ via iterated integration. Starting from the series definition, the multiple polylogarithm function can be expressed as an iterated integral as follows:
@@
If $(k_{1},\dotsc,k_{r})\in\mathbb{N}^{r}$ and $x_{1},\ldots,x_{r}$ are $N$th roots of unity, then (2.11) become the cyclotomic multiple zeta values of level $N$ which converges if $(k_{1},x_{1})\neq(1,1)$ (see [37] and [40, Ch. 15]). The level two cyclotomic multiple zeta values are called alternating multiple zeta values. When $x=1$ in (2.2), the series on the left-hand side was shown in [32, Theorem 2.2] to be expressible as a combination of alternating multiple zeta values.
@@
In particular, setting $x=\pm 1$ in Corollary 2.6 yields the two known results [34, Theorems 3.4 and 3.5]. As two examples, setting $q=1,2$ in (2.6) or letting $a=1/2$ in (2.16) and (2.5) yield
@@
Let $x=e^{i\theta}$ and $\theta\in(0,2\pi)$. For $a\in\mathbb{C}\setminus\mathbb{Z}$ and $m\in\mathbb{N}_{0}$, we have
@@
Proof.
From [33, P. 123, Example 6.12], we have
$\displaystyle\sum_{n=-\infty}^{\infty}\displaystyle\frac{e^{in\theta}}{n-x}$ $=2\pi i\displaystyle\frac{e^{i\theta x}}{1-e^{2\pi ix}}=-\pi\displaystyle\frac{e^{ix(\theta-\pi)}}{\sin(\pi x)}$
$=-\pi\cot(\pi a)x^{a}+\pi x^{a}i.$
On the other hand, noting the fact that
$\displaystyle\sum_{n=-\infty}^{\infty}\displaystyle\frac{x^{n}}{n-a}$ $=\displaystyle\sum_{n=1}^{\infty}\displaystyle\frac{x^{n}}{n-a}+\displaystyle\sum_{n=-\infty}^{0}\displaystyle\frac{x^{n}}{n-a}=\displaystyle\sum_{n=1}^{\infty}\displaystyle\frac{x^{n}}{n-a}-\displaystyle\sum_{n=1}^{\infty}\displaystyle\frac{x^{1-n}}{n+a-1}$
$=\operatorname{Li}_{1}(x;1-a)-x\operatorname{Li}_{1}(x^{-1};a).$
Hence, we obtain
$\operatorname{Li}_{1}(x;1-a)-x\operatorname{Li}_{1}(x^{-1};a)=-\pi\cot(\pi a)x^{a}+\pi x^{a}i.$
Finally, differentiating the above equation $m$ times with respect to a yields the desired result. ∎
An explicit formula for arbitrary high-order derivatives of $\cot(\pi a)$ can be found in [36, Thm. 2.3]. As an example, setting $m=1$ in (2.21) gives
@@
Due to the complexity of the higher-order derivative formulas for $\cot(\pi a)$, this paper retains the original form of the left-hand side of (2.21) rather than rewriting it into its right-hand equivalent. It should be noted that Goncharov [21, p.12] introduced the following multi-series version of the series appearing in the proof of Theorem 2.7:
$B(\varphi_{1},\ldots,\varphi_{m}|t_{1},\ldots,t_{m}):=\displaystyle\sum_{-\infty<k_{1}<\cdots<k_{m}<\infty}\displaystyle\frac{e^{2\pi i(\varphi_{1}k_{1}+\cdots+\varphi_{m}k_{m})}}{(k_{1}-t_{1})\cdots(k_{m}-t_{m})},$
where $\varphi_{i}\in(0,1)$. By investigating this function or similar generalizations, certain symmetry and parity results for multiple zeta values and their related variants can be derived, as seen in [20, 12].
@@ 3 Preliminaries
see [16, Section 3.3] and [28, Section 2.8]. From [35, Eqs. (2.5) and (2.9)], it is known that
@@
([17, Lemma 2.1]) Let $\xi(s)$ be a kernel function and let $r(s)$ be a rational function which is $O(s^{-2})$ at infinity. Then
@@
([32, Lemma 3.2]) For $|s|<1$, the following identities hold:
@@
([32, Lemma 3.3]) For nonnegative integer $n$, when $s\to-n$, we have
@@
([29, Eq. (2.6)]) Let $x$ be root of unity. For $n\in\mathbb{Z}$,
@@
The remainder of the proof follows a similar line to that of Theorem 2.1, and we therefore leave the detailed procedures to the interested reader. ∎
Obviously, letting $b=a$ in Theorem 4.1 yields Theorem 2.1. All results in this paper have been numerically verified by the authors using Mathematica, thereby ensuring their correctness. From Theorem 4.1, many other conclusions can also be derived. For example, taking $q=1$ and $b\rightarrow a+1/2$, and noting that $\lim_{x\rightarrow-1}D_{0}(x)=0$, we obtain
@@ 5 Alternative Expression for a Class of Euler-Apéry-like Series
In this section, we employ a method analogous to that in [34, Theorem 3.8] to derive an alternative expression of the theorem itself. We begin by presenting an introduction to Fuss-Catalan numbers and the relevant properties of their generating function.
@@
and satisfy the generating function relation [19, (2.9)]
@@
In [34, Section 2.2], the authors established that the function $G_{m}(x)$ is monotonically increasing on the interval $(m-1)^{m-1}/{m^{m}},(m-1)^{m-1}/{m^{m}}]$ and satisfies $0<G_{m}(x)\leq m/(m-1)\ (m\geq 2)$. In fact, from the generating function relation (5.40), it is straightforward to observe that $G_{m}>1/2$. This follows from a contradiction argument: if we assume $0<G\leq 1/2$ for $x\in[-(m-1)^{m-1}/{m^{m}},0)$, the identity (5.40) $G_{m}(x)=1+xG_{m}(x)^{m}$ would imply $G_{m}(x)>1/2$, which contradicts the assumption. Therefore, for any integer $m>1$, the function $G_{m}(x)$ takes values in the interval $(1/2,2]$.
@@
([34, Theorem 3.8]) For positive integers $m$ and $p$,
@@
Next, we employ a method analogous to the proof of [34, Theorem 3.8] to derive an alternative expression for (5.1).
@@
$\log^{k}(1-x)=(-1)^{k}k!\displaystyle\sum\limits_{n=1}^{\infty}\displaystyle\frac{{{x^{n}}}}{n}{\zeta_{n-1}}(\{1\}_{k-1})\quad(x\in[-1,1))$
and

### v2

# Contour Integration and Cyclotomic Apéry-Like Series Involving Generalized Binomial Coefficients

@@ 1 Introduction
The main aim of this paper is to explore the following cyclotomic Apéry-like series, which incorporate generalized binomial coefficients and are termed cyclotomic parametric Apéry-like series, using the method of contour integration:
@@
where $a\in\mathbb{C}\setminus\mathbb{Z},\ q\in\mathbb{N}$ and $z$ is an arbitrary root of unity with $(q,z)\neq(1,1)$. Here the generalized binomial coefficient $\binom{a}{b}$ is denoted by
@@
In particular, when $a$ is a half-integer in (1.2) and $2n+2a$ is a non-positive integer, the reciprocal of the parametric central binomial coefficient, $\binom{2n+2a}{n+a}^{-1}$, is interpreted as zero by convention.
@@ 2 Main Theorems and Corollaries
The second cyclotomic Apéry-like series in equation (2.1) above can be expressed in terms of multiple polylogarithms and $\log(2)$, as derived from the following proposition (see [30, Theorem 3.8]).
@@
In [30, Theorem 3.8], the formula omits two absolute value symbol; specifically $\log^{p-1-k}(x)$ should be written as $\log^{p-1-k}(|x|)$, and $\log^{k-j-l}(1-G_{m}(x)^{-1})$ should be written as $\log^{k-j-l}(|1-G_{m}^{-1}(x)|)$. This adjustment is necessary because it can be proven that for $x\in[-(m-1)^{m-1}/m^{m},(m-1)^{m-1}/m^{m}]$, the value of $G_{m}(x)\ (m\geq 2)$ lies within the interval $(1/2,2]$. In Section 5, we will employ a method similar to that used in the proof of [30, Theorem 3.8] to derive an alternative expression for equation (2.2). This approach will allow us to establish several identities involving multiple polylogarithm functions. Certainly, a direct integration approach should yield an alternative expression for equation (2.2). As can be readily observed from (2.9), the series on the left-hand side of (2.2) can be transformed into the following integral:
@@
Based on this integral, following a process similar to the proof of [29, Theorem 2.2], the desired result should be obtainable.
@@
which converges if $|x_{1}\cdots x_{j}|<1$ for all $j=1,\dotsc,r$. In particular, if $x_{1}=x,x_{2}=\cdots=x_{r}=1$, then $\operatorname{Li}_{k_{1},\ldots,k_{r}}(x,\{1\}_{r-1})$ is the single-variable multiple polylogarithm function $\operatorname{Li}_{k_{1},\ldots,k_{r}}(x)$. Clearly, $\operatorname{Li}_{1}(x)=-\log(1-x)\ (x\in[-1,1))$. The (2.11) can be analytically continued to a multi-valued meromorphic function on $\mathbb{C}^{r}$ (see [21, 35, 36]). In [20, Lemma 2.2], Kaneko and Tsumura provided a detailed explanation of how to analytically continue the single-variable multiple polylogarithm to the domain $\mathbb{C}\setminus[1,+\infty)$ via iterated integration. Starting from the series definition, the multiple polylogarithm function can be expressed as an iterated integral as follows:
@@
If $(k_{1},\dotsc,k_{r})\in\mathbb{N}^{r}$ and $x_{1},\ldots,x_{r}$ are $N$th roots of unity, then (2.11) become the cyclotomic multiple zeta values of level $N$ which converges if $(k_{1},x_{1})\neq(1,1)$ (see [34] and [37, Ch. 15]). The level two cyclotomic multiple zeta values are called alternating multiple zeta values. When $x=1$ in (2.2), the series on the left-hand side was shown in [29, Theorem 2.2] to be expressible as a combination of alternating multiple zeta values.
@@
In particular, setting $x=\pm 1$ in Corollary 2.6 yields the two known results [30, Theorems 3.4 and 3.5]. As two examples, setting $q=1,2$ in (2.6) or letting $a=1/2$ in (2.16) and (2.5) yield
@@
([33, Thm. 2.8]) Let $x=e^{i\theta}$ and $\theta\in(0,2\pi)$. For $a\in\mathbb{C}\setminus\mathbb{Z}$ and $m\in\mathbb{N}_{0}$, we have
@@
An explicit formula for arbitrary high-order derivatives of $\cot(\pi a)$ can be found in [32, Thm. 2.3]. As an example, setting $m=1$ in (2.21) gives
@@
Due to the complexity of the higher-order derivative formulas for $\cot(\pi a)$, this paper retains the original form of the left-hand side of (2.21) rather than rewriting it into its right-hand equivalent.
@@ 3 Preliminaries
see [15, Section 3.3] and [25, Section 2.8]. From [31, Eqs. (2.5) and (2.9)], it is known that
@@
([16, Lemma 2.1]) Let $\xi(s)$ be a kernel function and let $r(s)$ be a rational function which is $O(s^{-2})$ at infinity. Then
@@
([29, Lemma 3.2]) For $|s|<1$, the following identities hold:
@@
([29, Lemma 3.3]) For nonnegative integer $n$, when $s\to-n$, we have
@@
([26, Eq. (2.6)]) Let $x$ be root of unity. For $n\in\mathbb{Z}$,
@@
The above two power series expansions can be directly obtained using Faà di Bruno’s formula (see [15, Sec. 3.4]).
Let $H(s)$ be defined as the integrand in the integral in (4.33). It is not difficult to observe that the function $H(s)$ has the following poles in the complex plane: aba-b is a pole of order $q$, $n+a$ (where $n\in\mathbb{Z}$) are simple poles, and n-n (where $n\in\mathbb{N}$) are also simple poles.
From Lemma 3.5 and the two expansions above, we can provide the Laurent series expansion of $H(s)$ at s=abs=a-b as follows:
$H(s)$ $=4^{s}\displaystyle\sum_{k_{1},k_{2},k_{3},k_{4}=0}^{\infty}\displaystyle\frac{2^{k_{3}}C_{k_{1}}(a-b)C_{k_{2}}(a-b)D_{k_{3}}(2a-2b)}{k_{1}!k_{2}!k_{3}!}$
$\qquad\qquad\times\left((-1)^{k_{4}}\operatorname{Li}_{k_{4}+1}(x;1-b)-x\operatorname{Li}_{k_{4}+1}\Big(x^{-1};b\Big)\right)(s+b-a)^{k_{1}+k_{2}+k_{3}+k_{4}-q}.$
Therefore, the residue at s=abs=a-b can be obtained through direct computation as follows:
${\rm Res}[H(s),a-b]$ $=\displaystyle\frac{1}{(q-1)!}\displaystyle\lim_{s\rightarrow 0}\displaystyle\frac{d^{q-1}}{ds^{q-1}}\left((s+b-a)^{q}\Phi(s-a;x)\displaystyle\frac{\Gamma^{2}(1+s)}{\Gamma(1+2s)}\displaystyle\frac{4^{s}}{(s+b-a)^{q}}\right)$
$\qquad\qquad\times\left((-1)^{k_{5}}\operatorname{Li}_{k_{5}+1}(x;1-b)-x\operatorname{Li}_{k_{5}+1}(x^{-1};b)\right).$
Additionally, the residues at the simple poles $n+a$ ($n\in\mathbb{Z}$) and n-n ($n\in\mathbb{N}$) can be directly calculated as follows:
${\rm Res}[H(s),n+a]$ $=\displaystyle\lim_{s\rightarrow n+a}(s-n-a)\Phi(s-a;x)\displaystyle\frac{\Gamma^{2}(1+s)}{\Gamma(1+2s)}\displaystyle\frac{4^{s}}{(s+b-a)^{q}}$
$=x^{-n}\displaystyle\frac{4^{n+a}}{(n+b)^{q}\binom{2n+2a}{n+a}}$
and
${\rm Res}[H(s),-n]=(-1)^{q+1}\displaystyle\frac{n4^{-n}}{(n+a-b)^{q}}\binom{2n}{n}x^{n}\left(\operatorname{Li}_{1}(x;1-a)-x\operatorname{Li}_{1}(x^{-1};a)\right).$
Applying Lemma 3.1 yields
$\displaystyle\sum_{n=-\infty}^{\infty}{\rm Res}[H(s),n+a]+\displaystyle\sum_{n=1}^{\infty}{\rm Res}[H(s),-n]+{\rm Res}[H(s),a-b]=0.$
Summing these three contributions gives the statement of the Theorem 2.1. ∎
Obviously, letting $b=a$ in Theorem 4.1 yields Theorem 2.1. From Theorem 4.1, many other conclusions can also be derived. For example, taking $q=1$ and $b\rightarrow a+1/2$, and noting the fact that $\lim_{x\rightarrow-1}D_{0}(x)=0$, we obtain
@@ 5 Alternative Expression for a Class of Euler-Apéry-like Series
In this section, we employ a method analogous to that in [30, Theorem 3.8] to derive an alternative expression of the theorem itself. We begin by presenting an introduction to Fuss-Catalan numbers and the relevant properties of their generating function.
@@
and satisfy the generating function relation [18, (2.9)]
@@
In [30, Section 2.2], the authors established that the function $G_{m}(x)$ is monotonically increasing on the interval $(m-1)^{m-1}/{m^{m}},(m-1)^{m-1}/{m^{m}}]$ and satisfies $0<G_{m}(x)\leq m/(m-1)\ (m\geq 2)$. In fact, from the generating function relation (5.40), it is straightforward to observe that $G_{m}>1/2$. This follows from a contradiction argument: if we assume $0<G\leq 1/2$ for $x\in[-(m-1)^{m-1}/{m^{m}},0)$, the identity (5.40) $G_{m}(x)=1+xG_{m}(x)^{m}$ would imply $G_{m}(x)>1/2$, which contradicts the assumption. Therefore, for any integer $m>1$, the function $G_{m}(x)$ takes values in the interval $(1/2,2]$.
@@
([30, Theorem 3.8]) For positive integers $m$ and $p$,
@@
Next, we employ a method analogous to the proof of [30, Theorem 3.8] to derive an alternative expression for (5.1).
@@
$\log^{k}(1-x)=(-1)^{k}k!\displaystyle\sum\limits_{n=1}^{\infty}\displaystyle\frac{{{x^{n}}}}{n}{\zeta_{n-1}}(\{1\}_{k-1})\quad(x\in[-1,1)),$